### v1

# Tropicalization through the lens of Łukasiewicz logic, with a topos theoretic perspective

Tropicalization through the lens of Łukasiewicz logic, with a topos theoretic perspective
A. Di Nola, B. Gerla, G. Lenzi
1 Introduction
The main aim of this paper is to show that the topics of Łukasiewicz logic, semirings and tropical structures fruitfully meet, by combining the ideas of [10], [17], [38] and [14]. This gives rise to a topos theoretic perspective to Łukasiewicz logic, see [12]. Our aim seems to be completely in line with the spirit of the following remark from [34]:
The confluence of geometric, combinatorial and
logical-algebraic techniques on a common problem
is one of the manifestations of the unity of mathematics.”
Tropicalization, originally, is a method aimed at simplifying algebraic geometry and it is used in many applications. It can be seen as a part of idempotent algebraic geometry. Tropical geometry produces objects which are combinatorially similar to algebraic varieties, but piecewise linear. That is, tropicalization attaches a polyhedral complex to an algebraic variety, obtaining a kind of algebraic geometry over idempotent semifields.
Originally the idea was focused on complex algebraic geometry and started from a field (or ring) $K$ and its polynomials, an idempotent semiring $(S,\wedge,+,0,1)$ and a valuation $v:K\to S$. A semiring is an abelian monoid $(A,+,0)$ with a multiplicative monoid structure $(A,\cdot,1)$ satisfying the distributive laws and such that $a\cdot 0=0\cdot a=0$ for all $a\in A$. A semiring is idempotent if $+$ is idempotent.
Usually $S=G\cup\{\infty\}$ where $G$ is a totally ordered abelian group and $\infty$ is an infinite extra element, but we suggest to be more liberally inspired by [38]: first, $G$ can be any lattice ordered abelian group, and $S=G\cup\{\infty\}$ is what we call the Rump semifield of $G$ Moreover, we point out that the passage from $G$ to $S$ and conversely is functorial, even an equivalence. So we propose here a functorial tropicalization.
A polynomial $p$ with coefficients in $K$ is replaced by a polynomial $Trop(p)$ with coefficients in $S$. In the tropicalization $Trop(p)$ of $p$, the polynomial product of $K$ becomes $+$ in $S$, the polynomial sum in $K$ becomes the infimum in $S$. The constants $k\in K$ become their value $v(k)$. The idea is that the tropicalized polynomial $Trop(p)$ should be simpler than $p$ but retain the combinatorial structure of $p$.
In this paper we intend to contribute to what we mean by tropical mathematics, which for us, in a very broad sense, is the use of idempotent semirings in mathematics.
The most used semiring in tropicalization is the tropical semiring
$\mathbb{R}^{maxplus}=(\mathbb{R}\cup\{-\infty\},max,+,-\infty,0)$
(or its dual with min instead of max). It is used in many contexts, from algebraic geometry to optimization to phylogenesis. An application to differential equations is given in [22]. The variety generated by $\mathbb{R}^{maxplus}$ includes the semiring reduct of $[0,1]$, but not the other way round. Instead, if we consider the negative cone of the reals, $\mathbb{R}^{neg}=(\mathbb{R}^{\leq 0}\cup\{-\infty\},max,+,-\infty,0)$, we have that the varieties generated by $[0,1]$ and $\mathbb{R}^{neg}$ are the same. Another famous semiring is the tropical semiring $\mathbb{N}^{trop}=(\mathbb{N}\cup\{+\infty\},min,+,+\infty,0)$. Once again, the varieties generated by $\mathbb{N}^{trop}$ and $[0,1]$ are the same. However, $\mathbb{N}^{trop}$ and $[0,1]$ have different first order theories. Thus, looking at these two varieties from a logical point of view, some differences emerge. This justifies approaching tropical algebraic structures from a logical point of view.
One can try to apply the same tropicalization idea to more general frameworks, like universal algebraic geometry by Plotkin [35]. This paper proposes a possible generalization of tropicalization of algebraic structures using many valued logic and especially fuzzy logic. In fact, the algebraic structures of fuzzy Łukasiewicz logic (MV-algebras) have a natural structure of idempotent semiring. Note that MV-semirings are mentioned in [20] in the context of weighted logics.
Łukasiewicz logic is a fuzzy logic, that is a logic where the set of values is the real interval $[0,1]$, rather than the case $\{0,1\}$ of classical logic. The connectives of Łukasiewicz logic are $x\oplus y=min(x+y,1)$ (replacing OR) and $\neg x=1-x$ (replacing NOT). Then AND gets replaced by $x\odot y=max(0,x+y-1)$ (the Łukasiewicz product). We have the tertium non datur, $x\oplus\neg x=1$, the non contradiction law $x\odot\neg x=0$, but not the idempotency: $x\oplus x\not=x$. Despite the lack of idempotency, MV-algebras are deeply connected with idempotent structures.
Łukasiewicz logic can be axiomatized as follows, where $x\to y$ means $\neg x\oplus y$:
  1. 1.
    $x\to(y\to x)$
  2. 2.
    $(x\to y)\to((y\to z)\to(x\to z))$
  3. 3.
    $((x\to y)\to y))\to((y\to x)\to x)$
  4. 4.
    ($(\neg x\to\neg y)\to(y\to x)$.
The only rule is modus ponens: from $x\to y$ and $x$ derive $y$.
Like the semantic counterpart of classical logic is given by Boolean algebras, the semantic counterpart of Łukasiewicz logic is given by MV-algebras. These are algebraic structures generalizing Boolean algebras and widely used in many applications, from quantum mechanics to games to functional analysis. The interplay between semirings and MV-algebras is very interesting. Every MV-algebra is in a natural way (better, in two natural dual ways) a semiring, as explained in [15], [14].
Namely, MV-algebras are structures $(A,\oplus,0,1,\neg)$ where $(A,\oplus,0)$ is a commutative monoid, $\neg 0=1$, $x\oplus 1=1$, $\neg\neg x=x$ and ¬($\neg(\neg x\oplus y)\oplus y)=\neg(\neg y\oplus x)\oplus x)$.
MV-algebras are also lattices under the ordering $x\leq y$ such that there is $z$ with $y=x\oplus z$. We let also $x\odot y=\neg(\neg x\oplus\neg y)$ and $x\ominus y=x\odot\neg y=\neg(\neg x\oplus y)$.
The main example of MV-algebra is $[0,1]$ with $x\oplus y=min(x+y,1)$ and $\neg x=1-x$. By [13] it follows that $[0,1]$ generates the variety of MV-algebras.
[…]
In order to define other interesting MV-algebras it is convienient to introduce Mundici functor $\Gamma$, see [31]. $\Gamma$ is a functorial equivalence between MV-algebras and Abelian $\ell$-groups with strong unit. This equivalence was originally used in order to clarify the relations between $AF$ $C^{*}$-algebras and their $K_{0}$ group. $\Gamma(G,u)$ is the interval $[0,u]$ of $G$ where $x\oplus y=(x+y)\wedge u$ and $\neg x=u-x$.
Another important example is the Chang MV-algebra $C=\Gamma(Z\,lex\,$ $Z,$ (1,(1,( 1 , 0))0))0 ) ) where $lex$ denotes lexicographic product of groups. We call $V(C)$ the variety generated by $C$ and $V(C)$-algebras the elements of $V(C)$. The algebra $C$ has a very special role in MV-algebras theory. $\Gamma^{-1}(C)$ is the $K_{0}$ of Behncke-Leptin $C^{*}$-algebra $A_{0,1},$ see [32].
An equivalence related to $\Gamma$ is the one called $\Delta$ between abelian $\ell$-groups and perfect MV-algebras, see ([17]), that is, MV-algebras generated by their infinitesimals. Namely, if $G$ is an abelian $\ell$-group, then $\Delta(G)=\Gamma(Z\ lex\ G,(1,0))$. So, $C$ can also be defined as $\Delta(Z)$.
We can think of the logic of perfect MV-algebras as the logic of the quasi true or quasi false.
We focus on algebraic models of an equational extension of $\L$ukasiewicz logic. Actually we consider the extension of $\L$ukasiewicz logic given by the equation in the section 3.1.
It is often stressed that enriched valuations represent ”perturbations” of valuations of classical models. In our case we can speak of ”infinitesimal perturbation” of classical propositional logic, that in turn, means infinitesimal perturbation of Boolean algebras. Although it seems rather exotic to consider infinitesimally perturbed Boolean algebras as models of a propositional logic [33], it happens that these models can have an elegant algebraic characterization. Indeed, we can see that, albeit by means of a categorical equivalence, such models can be described by a Stone space and a family of lattice ordered abelian groups. Actually we will consider weak boolean products of lattice ordered abelian groups. It will be quite interesting that such a logic may have theories interpreted into points of the presheaf topos over the multiplicative monoid of integers.
The rest of the paper is organized as follows:
In section $2$ we discuss the relations between MV-algebras, semirings and semifields, showing that the category of MV-algebras is equivalent to semiring based categories.
In section $3$ we use the variety $V(C)$ of MV algebras generated by $C$ to give a functorial notion of tropicalization of an algebraic structure. We also present a geometric invariant $\theta$ for algebras from the variety $V(C)$ which has role in the definition of tropicalization functor.
In section $4$ we analyze the algebraic aspects of $\theta$ invariants, showing that they can be endowed with a rich algebraic structure. Such structures form a class of extended semiring (lattice-Bisemirings) and a new kind of two sorted first order structures called Coupled $\ell$-Bisemirings. The study of such structures could allow a classification of $\theta$ invariants.
In section $5$ we pass to the particular case of perfect MV-algebras, exploring the role of $\theta$ functor in moving from $V(C)$-algebraic varieties into tropical algebraic varieties.
In section $6$ we will show how a logic admitting only truth values infinitesimally close to boolean values, surprisingly, it can exhibit models that are functorially connected to a category of points of a non-commutative geometry.
In section $7$ we sketch some conclusions.
2 MV-algebras, semirings and semifields
So far we mentioned the role of semirings and idempotent semifields in the tropicalization issues. Now we observe that studies concerning the relationship between semirings and MV-algebras started in [14] and developed in subsequent papers, see ([15],[18], [19], [21]). Furthermore, via semirings, in ([41]) Łukasiewicz logic is connected with Weighted Logics, introduced by Droste and Gastin ([20]). A generalisation of MV-semirings, actually the Near semirings, are used in [4] to represent a class of quantum structures.
Recall that in every additively idempotent semiring $(S,+,\cdot,0,1)$ there is a natural partial order such that $s\leq t$ if and only if $s+t=t$. Homomorphisms of additively idempotent semirings are increasing.
Definition 2.1.
An MV-semiring $(A,+,\odot,0,1)$ is a commutative idempotent semiring such that for every $s\in A$ there is a greatest element $s^{*}$ such that $s\odot s^{*}=0$ and we have $s+t=(s^{*}\odot(s^{*}\odot t)^{*})^{*}$ for every $s,t\in A$.
We call homomorphism of MV-semirings a homomorphism of semirings preserving *. In fact, not every homomorphism of semirings preserves *.
For instance, take the MV-algebras $\{0,1/2,1\}$ and $\{0,1\}$, take the map $h:\{0,1/2,1\}\to\{0,1\}$ such that $h(0)=0$, $h(1/2)=0$, $h(1)=1$. Then $h$ is a homomorphism of semirings, but does not preserve *, since $h((1/2^{*}))=h(1/2)=0$ and $h(1/2)^{*}=0^{*}=1$.
Definition 2.2.
A coupled semiring is a structure $(A,+,\odot,0,1,+^{\prime},\odot^{\prime},\alpha)$ such that $(A,+,\odot,0,1)$ and $(A,+^{\prime},\odot^{\prime},1,0)$ are commutative idempotent semirings, $\alpha$ is an isomorphism between them with $\alpha(\alpha(x))=x$, and $x+y=x\odot^{\prime}(\alpha(x)\odot y)$.
A morphism of coupled semirings is defined as a map which preserves all the operations of the structure. The idea of coupled semiring is extended in [9] to residuated lattices satisfying the law of double negation.
We consider the categories of MV-algebras, MV-semirings and coupled semirings, with their corresponding homomorphisms. In [2] it is shown that such categories are equivalent, we repeat here the main steps of the proof.
Theorem 2.1.
The categories of MV-algebras and MV-semirings are equivalent.
Proof.
We define a functor $S$ from MV-algebras to MV-semirings. Given an MV-algebra $A$, let $S(A)$ be the MV-semiring associated to $A$. Given a homomorphism $h:A\to B$ between two MV-algebras, then $h$ preserves $+$, $\odot$ and , so $h:S(A)\to SB$ is a homomorphism of MV-semirings and we let $Sh=h$.
Suppose $k:SA\to SB$ is a morphism of MV-semirings associated to MV-algebras $A$ and $B$. Then $k$ preserves $\odot$ and , so $k$ preserves also $\oplus$. Hence, $k$ preserves the MV-algebra structure of $A$ and $k$ is a homomorphism of MV-algebras, and $Sk=k$. Clearly if $Sk=Sk^{\prime}$ then $k=Sk=Sk^{\prime}=k^{\prime}$. So $S$ is full and faithful
Finally, every MV-semiring $(S,\odot)$ is associated to the MV-algebra $(S,0,\oplus,*)$ where $a\oplus b=(a^{*}\odot b^{*})^{*}$, so $S$ is surjective on objects.
Summing up, $S$ is an equivalence ([30]). ∎
In the same way one proves:
Corollary 2.1.
The categories of MV-algebras and coupled semirings are equivalent.
Consider now the subcategory of MV-algebras given by Boolean algebras. A Boolean algebra is an MV-algebra where $x^{2}=x$, where we use the notation $x^{2}=x\odot x$. So by restricting the above results we have:
Corollary 2.2.
The categories of Boolean algebras and MV-semirings verifying $x^{2}=x$ are equivalent. The categories of Boolean algebras and coupled semirings verifying $x^{2}=x$ are equivalent.
Now let us turn to semifields, which are our tropical objects.
A semifield is a semiring $(F,+,\times,0,1)$ where the multiplication is commutative and every nonzero element has an inverse with respect to the multiplication. A semifield is idempotent if $x+x=x$ for every $x\in F$.
We have several categorial results relating MV-algebras and semifields. First of all we have:
Theorem 2.2.
[38] There is an equivalence $\rho$ between the category of idempotent semifields and the category of abelian $\ell$-groups.
The equivalence in Theorem 2.2 acts by associating with every idempotent semifield $(F,+,\times,0,1)$ the abelian $\ell$-group $(F\setminus\{0\},\cdot,1)$..
The equivalence $\rho$ can be refined as follows. Let $(F,+,\times,0,1)$ be an idempotent semifield. Let $u\in F$. We say that $u$ is a strong unit of $F$ is $u\not=0$ and for every $x\in F,x\not=0$, there is $n$ such that $x^{n}\leq u$. Again, by removing the zero of the semifield we get the following:
Proposition 2.1.
There is an equivalence between idempotent semifields with a strong unit and abelian $\ell$-groups with a strong unit.
Now by composing $\Gamma$ and $\rho$ we obtain:
Theorem 2.3.
([19]) There is an equivalence between idempotent semifields with a strong unit and MV-algebras.
The previous theorem supports the claim that tropical geometry can be made not only on semifields, but even on MV-algebras. That is we get a perspective for a tropical geometry of algebraic models of $\L$ukasiewicz logics.
Moreover, note that Theorem 2.3 specializes to subvarieties as follows:
Corollary 2.3.
For every subvariety $V$ of MV-algebras there is an equivalence between $V$ and a category of semifields with strong unit definable in first order logic over the language of semifields with strong unit.
Proof.
Every variety $V$ of MV-algebras is defined by a finite set $\Sigma_{V}$ of equations in one variable. Now for every abelian $\ell$-group $G$, the condition that $\Gamma(G,u)$ satisfies $\Sigma_{V}$ is definable in $G$ by a first order formula $\phi$ over the $\ell u$-group language of $G$. In turn this condition is equivalent to a first order formula $\phi^{nonz}(x)$ over $\rho^{-1}(G)$ in the language of semifields. In fact, $\phi^{nonz}(x)$ is obtained by restricting the quantifiers of $\phi$ to the nonzero elements of the semifield. ∎
By composing the functors $\Delta$ and $\rho$ we have another result concerning perfect MV-algebras (which are not a variety):
Theorem 2.4.
There is an equivalence between the categories of idempotent semifields and perfect MV-algebras.
Theorems 2.3 and 2.4 are the first steps of a comparison between tropical geometry and geometry over MV-algebras.
3 Rump tropicalization and its generalization via $V(C)$ algebras
Our aim is to show how much the topics of $V(C)$-algebras, their universal algebraic geometry, logic, semirings and Tropical structures fruitfully meet. We will show how a special class of semirings, which are reducts of MV-algebras, is related with issues of Dequantization (see [38]), Tropical Geometry and Non Commutative Geometry. This will be seen through the lens of Super Łukasiewicz Logic, [16], obtaining a logical and geometric view of tropical structures.
We propose a functor $\theta$ from the category $V(C)$ to $Set$ which, in our opinion, plays the role of a tropicalization of algebraic structures and extends Rump tropicalization of abelian $\ell$-groups, see [38]. We recall that Rump associates to every abelian $\ell$-group a semifield such that:
  1. 1.
    the semifield is idempotent;
  2. 2.
    the correspondence is a categorical equivalence.
We can say that the tropicalization of a category (in the sense of Rump) is a functor satisfying the two properties above.
3.1 $\theta$ invariant of $V(C)$-algebras
In this section we will focus on the equational class of $V(C)$ which is a particularly natural extension of Boolean algebras inside the variety of MV-algebras. Indeed, as has been noted in [8], the MV-algebras in $V(C)$ are Boolean algebras “perturbed” with infinitesimals. More precisely, every element is infinitely close to a Boolean element. From the point of view of universal algebra, $V(C)$ is an atom in the lattice of subvarieties of the variety of MV-algebras, where the variety $Bool$ of Boolean algebras is the smallest subvariety.
Let us define infinitesimal an element $x$ such that $nx\leq\neg x$ for every positive integer $n$. The idea is that $x\leq 1/n$, but $1/n$ does not necessarily exist in an MV-algebra. The set of all infinitesimals of an MV-algebra $A$ is called the radical of $A$ and is denoted by $Rad(A)$. A perfect MV-algebra is an MV-algebra generated by its radical, hence by its infinitesimals.
Note that $[0,1]$ has no infinitesimals, whereas $C$ has them. Indeed, $C$ consists only of infinitesimals and negations of infinitesimals, that is, is perfect. Every MV-algebra has a largest perfect subalgebra, consisting of the radical and coradical.
The co-radical of the Lindenbaum algebra of first order Łukasiewicz logic is given by the formulas which are true but not provable, see [39]. It would be interesting to describe the radical of this MV-algebra.
MV-algebras form a variety having countably many subvarieties. We will focus on the subvariety generated by $C$, denoted by $V(C)$. This variety is axiomatized by
$(2x)^{2}=2(x^{2})$
where $2x$ means $x\oplus x$ and $x^{2}$ means $x\odot x$. We remark that the corresponding logic has some nice properties, for example, the propositional logic is complete with respect to the MV-chain $\Delta(R)$, and a first order logic is complete with respect to all Perfect MV-chains.
We note that the above identity is used in [11] in defining the variety of DL-algebras and its free algebras.
In $V(C)$, every element is infinitely close to a Boolean element, that is to an element $x=x\oplus x$. So, we can consider $V(C)$-algebras as perturbed Boolean algebras. Note that infinitesimals cannot exist in Boolean algebras, but they occur in mathematics, from nonstandard analysis [36] to first order fuzzy logic [39]. More generally, every MV-algebra $A$ has a Boolean part $B(A)$, a Boolean algebra consisting of the elements $x$ such that $x=x\oplus x$, and a perfect part $P(A)$, the perfect MV-algebra generated by the infinitesimals of $A$. If $A\in V(C)$, then $A$ is a combination of $P(A)$ and $B(A)$, a perfect MV-algebra and a Boolean algebra. But by the functor $\Delta$, a perfect MV-algebra corresponds to an Abelian $\ell$-group. So in a sense, an MV-algebra in V(C) is a combination of a Boolean algebra and an Abelian $\ell$-group. We can say that the Boolean algebra carries a qualitative information, and the $\ell$-group carries a quantitative information.
Definition 3.1.
Given any MV-algebra $A\in V(C)$, we let
$\theta(A)=\{x\in A|x\geq 2x^{2}\},$
$\theta^{*}(A)=\{x\in A|x\leq 2x^{2}\},$
where $x^{2}=x\odot x$ and $2x=x\oplus x$.
Since $\theta(A)$ is the zerosets of an MV-polynomial, that is an algebraic variety of the $V(C)$, it is invariant with respect to isomorphic MV-algebras. Note that for every $A\in V(C)$,
  • $B(A)\subseteq\theta(A)$ (in particular, $0,1\in\theta(A)$),
  • $Rad(A)\subseteq\theta(A)$
  • $co-Rad(A)\cap\theta(A)=\{1\}$.
Lemma 3.1.
Let $A$ be an MV-algebra in $V(C)$. Let $x\in A$. The following are equivalent.
  • $x\geq 2x^{2}$, that is, $x\in\theta(A)$.
  • $x$ is a sum of a boolean and an infinitesimal.
  • for every prime ideal $P$, $x\ mod\ P$ is infinitesimal or $1$.
Proof.
We first prove that the second condition implies the first. Since every algebra in $V(C)$ is subdirects product of linearly ordered $V(C)$-algebras, it is enough to consider linearly ordered $V(C)$-algebras. In this case, the boolean part of $x$ is 00 or $1$, so $x$ is either an infinitesimal or $1$. In both cases, we have $x\geq 2x^{2}$.
Conversely, we know that every element $x$ of an MV-algebra in $V(C)$ decomposes as $x=(b\oplus\epsilon)\ominus\tau$, where $b$ is boolean and $\epsilon,\tau$ infinitesimals such that $\epsilon\wedge\tau=0$. Since every algebra in $V(C)$ is subdirects product of linearly ordered $V(C)$-algebras $L_{i}$ for $i$ in an index set $I$, we can consider the $i$-th component of the expression: $x_{i}=(b_{i}\oplus\epsilon_{i})\ominus\tau_{i}$. If $b_{i}=0$ then we must have $\tau_{i}=0$. If $b_{i}=1$, then from the hypothesis we infer $x_{i}\geq 2x_{i}^{2}$ and this implies $\tau_{i}=0$. Summing up, $\tau=0$ and $x=b\oplus\epsilon$.
The first condition implies the third since it is essentially equational. Conversely, if the third condition holds, then $x\geq 2x^{2}$ holds modulo any prime $P$ of $A$, so it holds in $A$.
We note:
Lemma 3.2.
If $A\in V(C)$, then $\theta(A)$ is closed under ,,$\oplus$,$\odot$,$\wedge$,$\vee$ , ⊙ , ∧ , ∨.
Proof.
Suppose $x\geq 2x^{2}$ and $y\geq 2y^{2}$. Since every algebra in $V(C)$ is subdirect product of linearly ordered $V(C)$-algebras $L_{i}$ for $i$ in an index set $I$, we can consider the $i$-th component of $x,y$ and we have $x_{i}\geq 2x_{i}^{2}$ and $y_{i}\geq 2y_{i}^{2}$. Then $x_{i}$ and $y_{i}$ can only be infinitesimal or $1$. So all their combinations under ,,$\oplus$,$\odot$,$\vee$,$\wedge$ , ⊙ , ∨ , ∧ are also infinitesimals or $1$ and they belong to $\theta(L_{i})$. ∎
In general $\theta(A)$ is not closed under negation. In fact we have:
Lemma 3.3.
Let $A\in V(C)$. The following are equivalent.
  • $\theta(A)$ is closed under negation.
  • $A=\theta(A)$.
  • $\theta(A)$ is an MV-algebra.
  • $A$ is a Boolean algebra.
Proof.
If $A=\theta(A)$ then infinitesimals cannot exist in $A$, otherwise coinfinitesimals would exist in $\theta(A)$, but coinfinitesimals never satisfy $x\geq 2x^{2}$ unless $x=1$.
If $A$ is a Boolean algebra then $x=x^{2}=2x^{2}$ for every $x\in A$, hence $A=\theta(A)$. ∎
4 Semirings, $\theta$ reducts and $\ell$-Bisemirings
In this section we will analyze the algebraic aspects of $\theta$ invariant. We see that the algebraic variety $\theta(A)$ can be endowed with a rich algebraic structure. This opens the door to classifications of these varieties.
Definition 4.1.
An $\ell$-Bisemiring an algebra $(S,\wedge,\vee,\odot,\oplus,0,1)$ such that:
  1. 1.
    $(S,\wedge,\vee,0,1)$ is a bounded distributive lattice;
  2. 2.
    $(S,\wedge,0,1,\oplus)$ is an idempotent commutative semiring; $(S,\vee,0,1,\odot)$ idempotent commutative semiring.
Clearly, every MV-algebra is an $\ell$-Bisemiring.
Definition 4.2.
A Coupled $\ell$-Bisemirings is a two sorted first order structure $(S,T,\nu)$ where
  1. 1.
    $S$ and $T$ are $\ell$-Bisemirings;
  2. 2.
    $\nu$ is an isomorphism between $S$ and $T$ swapping $\vee$,$\wedge$ , ∧ and $\oplus$,$\odot$ , ⊙.
We denote by $C\ell BS$ the category of Coupled $\ell$-Bisemirings and their morphisms (as two sorted structures of first order logic). Namely, a morphism $(S,T,\nu)\to(S^{\prime},T^{\prime},\nu^{\prime})$ is a pair of $\ell$-Bisemiring homomorphisms $(f,g)$ such that $f:S\to S^{\prime}$, $g:T\to T^{\prime}$, and $\nu^{\prime}\circ f=g\circ\nu$.
In the previous section we introduced functors $\theta$ and $\theta^{*}$ from $V(C)$-algebras to $Set$. We observe that $\theta(A)$ and $\theta^{*}(A)$ are algebraic varieties in the sense of algebraic geometry, despite the structure used in defining them is the one of MV-algebras instead of rings or fields. The negation in $A$ is an anti isomorphism between $\theta(A)$ and $\theta^{*}(A)$.
Let us define $\Theta(A)=(\theta(A),\theta^{*}(A),\neg)$.
Theorem 4.1.
For every $A\in V(C)$, the structure $\Theta(A)$ belongs to $C\ell BS$. The map sending $A$ to $\Theta(A)$ is functor from $V(C)$ to $C\ell BS$.
Proof.
In fact, if $h:A\to B$ is a homomorphism of MV-algebras in $V(C)$, then $h$ sends $\theta(A)$ in $\theta(B)$ because $\theta(A)$ is the zeroset of an MV-algebraic polynomial, namely, $2x^{2}\ominus x$. So we can associate to $h$ the map $\theta(h):\theta(A)\to\theta(B)$ obtained by restricting $h$.
More precisely, the functor $\theta$ goes from the category $V(C)$ seen as a category of MV-algebras, to a category of structures $(S,\wedge,\vee,\oplus,\odot)$ with four binary operations (the first two are distributive lattice operations). Note that any homomorphism of MV-algebras preserves the four operators.
The same holds for $\theta^{*}$.
In particular, $(\theta^{*}(A),\wedge,\oplus,1,0)$ is again a commutative, additively idempotent semiring.
On the other hand, with every MV-algebra $A$ we can associate a Coupled $\ell$-Bisemiring by considering $(A,A,id_{A})$. Thus we have a functor from the category of MV-algebras to $C\ell BS$ and a different functor from the category of $V(C)$ to $C\ell BS$.
The structure $(\theta(A),\oplus,\odot,\vee,\wedge,0,1)$ is a Coupled $\ell$-Bisemiring: we want now to characterize such structures.
For $A\in V(C)$ we note that an element $x\in\theta(A)$ is boolean in $A$ if and only if $x^{2}=x$, and $x$ is infinitesimal in $A$ if and only if $x^{2}=0$. This means that booleans and infinitesimals are definable internally in $\theta(A)$, since multiplication is available in $\theta(A)$. We begin with a lemma which allows us to put together a boolean algebra and a perfect algebra.
Lemma 4.1.
Let $B$ be a boolean algebra and $P$ a perfect MV-algebra. There is an MV-algebra $A\in V(C)$ such that $B(A)=B$ and $Rad(A)=Rad(P)$.
Proof.
We perform the following detour. Let $B^{\prime}$ be a boolean algebra such that $B\subseteq B^{\prime}\times\{0,1\}$ ($B^{\prime}$ exists by Stone Theorem on Boolean algebras). Consider $A=\{(b,p)\in B^{\prime}\times P|(b,p\ mod\ Rad(P))\in B\}$.
Note that $A\in V(C)$. Moreover, $B(A)=B$ and $Rad(A)$ is the set of the elements $(0,p)$ with $p\in Rad(P)$. So $Rad(A)$ is isomorphic to $Rad(P)$.
Now we have the desired characterization of $\theta$ of MV-algebras in $V(C)$:
Lemma 4.2.
Let $(S,T,\nu)$ be a Coupled $\ell$-Bisemiring. Then $S=\theta(A)$ for some $A\in V(C)$ if and only if:
  • $(Inf(S),\oplus,\odot,\wedge,\vee)$ is the radical of an MV-algebra, where $Inf(S)$ is the set of $x\in S$ such that $x^{2}=0$.
  • $(BOOL(S),\oplus,\odot)$ is a Boolean algebra, where $\vee$=$\oplus$ = ⊕ and $\wedge$=$\odot$ = ⊙ and $Bool(S)$ is the set of $x$ such that $x^{2}=x$.
Proof.
$\theta(A)$ has the required properties for every $A\in V(C)$.
Conversely, to obtain $A$ from $S$ it is enough take an MV-algebra $A\in V(C)$ such that $B(A)=Bool(S)$ and $Rad(A)=Inf(S)$. One such MV-algebra is built in the previous lemma. ∎
By specializing the result to perfect MV-algebras we have:
Corollary 4.1.
Let $(S,T,\nu)$ be a Coupled $\ell$-Bisemiring. Then $S=\theta(A)$ for some perfect MV-algebra $A$ if and only if:
  • $(Inf(S),\oplus,\odot,\wedge,\vee)$ is the radical of an MV-algebra, where $Inf(S)$ is the set of $x\in S$ such that $x^{2}=0$.
  • $(Bool(S),\oplus,\odot)$ is a two element Boolean algebra, where $\vee$=$\oplus$ = ⊕ and $\wedge$=$\odot$ = ⊙ and $Bool(S)$ is the set of $x$ such that $x^{2}=x$.
It would be interesting to classify all MV-algebras with a given $\theta$. Probably one should consider actions of boolean algebras on monoids (a link with the arithmetic site of [12]?).
We have seen that the algebraic variety $\theta(A)$, with $A\in V(C)$, have a lot of algebraic structure. This is confirmed by the following operations, all available in $\theta(A)$:
  • $bool:A\to B(A)$, $bool(x)=2x^{2}$ (the boolean part of $x$)
  • $inf:A\to Rad(A)$, $inf(x)=x\wedge\neg x$ (the infinitesimal part of $x$)
  • $neg_{b}:A\to B(A)$, $\neg_{b}(x)=\neg(bool(x))$.
5 Tropicalization of Perfect MV-algebras
The aim of this section is to show how a generalization of the concept of tropicalization of algebraic structures makes it possible allows one to discover interesting inter-relationship between categories of algebraic models of $\L$ukasiewicz logics and their algebraic geometries through a tropicalization functor.
In this section we continue to describe the algebraic aspects of the algebraic varieties $\theta(A)$ in the cases where $A$ is a Perfect MV-algebra. Perfect MV-algebras are a full subcategory of $V(C)$. We remark that they Perfect MV-algebras do not form an equational class.
We explore the role of $\theta$ functor in composition with other functors acting in tropical geometry, and more, its role in moving from $V(C)$-algebraic varieties into tropical algebraic varieties.
Definition 5.1.
Let $TOP$ be the category of positive cones of $\ell$-groups with a top, where morphisms preserve the positive cone and the top. Let $BOT$ be the category of negative cones of $\ell$-groups with a bottom, where morphisms preserve the negative cone and the bottom. Let $PERF$ be the category of perfect MV-algebras.
Theorem 5.1.
$\theta$ extends to an equivalence between $PERF$ and $TOP$. $\theta^{*}$ extends to an equivalence between $PERF$ and $BOT$.
Proof.
Let $P$ be a perfect MV-algebra. Then $\theta(P)=Rad(P)\cup\{1\}$ and $Rad(P)$ is the positive cone of the $\ell$-group $\Delta^{-1}(P)$. Then it is easy to check that the claim of the theorem holds. ∎
Let $G$ be an abelian $\ell$-group. The $\Delta$ functor maps $G$ in a perfect MV-algebra $\Delta(G)$. Let $\ell\cal G$ be the category of abelian $\ell$-groups. Applying the functors $\theta,\theta^{*}$ we obtain:
Theorem 5.2.
$\theta\circ\Delta$ is a categorial equivalence from $\ell\cal G$ to $TOP$. $\theta^{*}\circ\Delta$ is a categorial equivalence from $\ell\cal G$ to $BOT$.
Now if $G$ is an abelian l-group, we define its tropicalization $Trop(G)$ as the additively idempotent semifield obtained by adding a point $\infty$- ∞ to $G$. The tropicalization can be considered as a functor from $\ell\cal G$ to the category of idempotent semirings. The inverse functor takes an additively idempotent semifield $S$ and maps it to the abelian l-group $Detrop(S)$ obtained by deleting the zero of $S$.
So, given an additively idempotent semifield $S$, we can define two functors $\mathbb{F},\mathbb{F^{*}}$ by a triple composition of functors
$\mathbb{F}(S)=\theta(\Delta(Detrop(S))),$
$\mathbb{F^{*}}(S)=\theta^{*}(\Delta(Detrop(S))),$
Theorem 5.3.
$\mathbb{F}$ is an equivalence from the category of additively idempotent semifields to $TOP$. $\mathbb{F^{*}}$ is an equivalence from the former category to $BOT$.
So the functor $\mathbb{F}$ is an equivalence, and gives a direct link between tropical structures ( additively idempotent semifields) and our $\theta$ structures over Perfect MV-algebras. The construction is functorial, and is more general than dequantization, see [38], in the sense that dequantization leads only to the construction of single structures like the (max,plus)-semiring.
So, the above theorem seems to be the answer to what we propose in the introduction, where, namely, we proposed to build up a functorial tropicalization of an algebraic structure. Indeed, we see the elements of $TOP$ as tropicalizations of perfect MV-algebras. It remains to tropicalize in a reasonable sense more general MV-algebras.
6 MV-algebras and the topos $\widehat{\mathbb{N}^{\times}}$
In this section, we will show how a logic admitting only truth values infinitesimally close to boolean values, can surprisingly exhibit models that are functorially connected to a category of points of a non-commutative geometry.
We are inspired by [12] where the authors apply topos theory and tropical geometry for a possible road to Riemann Hypothesis. The hope is that, as schemes opened the way for Fermat’s theorem, topoi open the way for Riemann hypothesis. In [12] we found the crucial construction presented here by
Definition 6.1.
We denote by $\mathbb{N}^{\times}$ the monoid of integers with multiplication and by $\widehat{\mathbb{N}^{\times}}$ its presheaf category of the functors from $\mathbb{N}^{\times}$, seen as a category, to $Set$.
The points of the topos $\widehat{\mathbb{N}^{\times}}$ form a category which is studied in [12]. We aim to establish some model theoretic properties of classes of perfect MV-algebras. Then, we will specialize the obtained results to the category of $Perf^{\mathbb{Q}}$, below defined, and proceeding along this road to get model theoretic properties of the class of points of $\widehat{\mathbb{N}^{\times}}$.
Indeed we aim to give a MV-algebraic characterization of the points of the presheaf topos $\widehat{\mathbb{N}^{\times}}$ and the logical consequences of this characterization.
We associate points of $\widehat{\mathbb{N}^{\times}}$ to theories in $L_{p}$, the extension of Lukasiewicz logic with the equation defining the variety $V(C)$.
Recall that by $\Delta$ we denote the categorical equivalence between abelian $\ell$-groups and perfect MV-algebras, see ([17]). Let us call $\Delta^{\mathbb{Q}}$ the restriction of the functor $\Delta$ to $Ab^{\mathbb{Q}}$, i.e., the category of abelian $\ell$-subgroups of $\mathbb{Q}$. Then we set $Perf^{\mathbb{Q}}:=\Delta^{\mathbb{Q}}(Ab^{\mathbb{Q}})$. Likewise we can consider the restriction $\theta^{\mathbb{Q}}$ of $\theta$ to the range of $\Delta^{\mathbb{Q}}$ and $TOP^{\mathbb{Q}}$ the image of $\Delta^{\mathbb{Q}}$.
Definition 6.2.
Call an ordered monoid subtractive if $x\leq y$ holds if and only if there is $z$ with $y=x+z$.
Let us call $Canc^{\mathbb{Q}}$ the category of cancellative subtractive lattice ordered abelian submonoids of $\mathbb{Q}^{+}$ with an extra bottom element.
The paper [8] gives an equivalence between abelian $\ell$-groups and cancellative subtractive lattice ordered abelian monoids with bottom element. The idea is to associate to an $\ell$-group its positive cone plus an infinite point. By restricting the equivalence to the subgroups of $\mathbb{Q}$, we obtain an equivalence between $Ab^{\mathbb{Q}}$ and $Canc^{\mathbb{Q}}$. Summing up we have:
Theorem 6.1.
The categories $Ab^{\mathbb{Q}}$, $TOP^{\mathbb{Q}}$, $Canc^{\mathbb{Q}}$ and $Perf^{\mathbb{Q}}$ are pairwise equivalent.
Now we give some translation lemmas. Note that if $A$ is an MV-algebra then $Rad(A)$ is closed with respect to operations $\oplus$ and $\ominus$.
Lemma 6.1.
Let $A$ be a perfect MV-algebra and let $\phi$ be a formula of the language of MV-algebras. Then $A\models\phi(x_{1},\ldots,x_{n})$ if and only if $Rad(A)\models Rad(\phi)(y_{1},\ldots,y_{n})$, where $Rad(\phi)$ is a formula in the language of $\oplus$,$\ominus$ , ⊖, and $y_{i}=x_{i}$ if $x_{i}$ has type Rad, and $y_{i}=\neg x_{i}$ otherwise.
Proof.
Since $A$ is perfect, $A$ is the union of the radical and the coradical. So we can suppose that the variables of $\phi$ have type $Rad$ or $Corad$.
The type of a term (rad or corad) is defined inductively as follows. 00 has type rad. $\neg t$ has type rad if and only if $t$ has type corad. Finally $t\oplus u$ has type rad if both $t,u$ have type rad, and corad otherwise.
The radical is equipped with the functions $\oplus$,$\ominus$ , ⊖.
We prove inductively that every term of type Rad is a combination of $\oplus$,$\ominus$ , ⊖ of variables of type rad and negated variables of type corad.
Suppose a term $t$ is of type Rad.
If $t$ is a variable then $t$ has the desired form.
If $t=\neg u$ and $u$ is a variable then $t$ has the desired form.
If $t=u\oplus v$ then $u,v$ have type rad and we are done.
If $t=\neg\neg u$ then $u$ has the desired form.
Suppose $t=\neg(u\oplus v)=\neg u\odot\neg v$. If $u,v$ have type rad then $t$ has type corad, which is a contradiction.
If $u,v$ have type corad then $t=0$.
If $u$ has type rad and $v$ has type corad then $\neg v$ has type rad and $t=(\neg v)\ominus u$ which has the desired form. Similarly we proceed if $u$ is corad and $v$ is rad.
We can suppose that every term in $\phi$ has type rad. In fact, suppose $t=u$ have different types. Then $t=u$ can be replaced by false. If $t=u$ occurs in $\phi$ and $t,u$ have type corad, then $t=u$ can be replaced by $\neg t=\neg u$. So the only case remaining is that only terms of type rad occur in $\phi$.
Now the formula $\exists x:corad.\phi(\neg x)$ can be replaced by $\exists x:rad.\phi(x)$. In this way $\phi$ is transformed in a formula whose variables are all of type rad, and all terms are combinations of $\oplus$ and $\ominus$ of variables.
Going from radicals to groups we get:
Lemma 6.2.
Take a formula $\phi$ in the language of $\oplus$,$\ominus$ , ⊖. There is a formula $\Delta_{0}(\phi)$ in the language of ordered groups, effectively built from $\phi$, such that for every lattice ordered abelian group $G$, we have $Rad(\Delta(G))\models\phi(x_{1},\ldots,x_{n})$ if and only if $G\models\Delta_{0}(\phi)(x_{1},\ldots,x_{n})$.
Proof.
This is because in $G$ we have $x\oplus y=x+y$, $x\ominus y=0\vee(x-y)$. ∎
Now by combining the translations $\Delta_{0}$ and $Rad$ we obtain:
Theorem 6.2.
Take a formula $\phi$ in the language of MV-algebras whose variables range over the radical or the coradical. There is a formula $\Delta(\phi)$ in the language of ordered groups, effectively built from $\phi$, such that for every lattice ordered abelian group $G$, we have $\Delta(G)\models\phi(x_{1},\ldots,x_{n})$ if and only if $G\models\Delta(\phi)(y_{1},\ldots,y_{n})$, where $y_{i}=x_{i}$ if $x_{i}$ ranges over the radical, and $y_{i}=\neg x_{i}$ otherwise.
If we specialize from $V(C)$ to $Perf_{\mathbb{Q}}$ we obtain more:
Corollary 6.1.
The map $\Theta_{\mathbb{Q}}$ sending $A\in Perf_{\mathbb{Q}}$ to $\Theta(A)$ is an equivalence of categories from $Perf_{\mathbb{Q}}$ to a full subcategory of $C\ell BS$.
Proof.
Let $\mathcal{R}$ be the full subcategory of $C\ell BS$ consisting of $\Theta(A)$ for $A\in Perf_{\mathbb{Q}}$. $\Theta_{\mathbb{Q}}$ is a full and faithful functor from $Perf_{\mathbb{Q}}$ to $\mathcal{R}$. ∎
We will see that the duality $\Delta$ can be used to transfer logical information between abelian $\ell$-groups and perfect MV-algebras, in particular, regarding first order logic. Something similar happens for the duality $\Gamma$, but not quite for first order logic, rather geometric logic in the sense of topos theory, since both MV-algebras and $\ell$ groups with strong units are axiomatized in geometric logic.
Here we recall that a class of models is geometric if it is axiomatized by geometric sequents in the sense of topos theory, see [6].
Geometric logic is a fragment of infinitary first order logic of a constructive flavor where one allows infinite disjunctions, finite conjunctions and existential quantifiers (no negation).
We start by investigating the first order theory of the non trivial subgroups of the rationals as ordered groups. Then via the $\Delta^{\mathbb{Q}}$ functor we can transfer the obtained results to the first order theory of $Perf^{\mathbb{Q}}$ and to points of such a natural topos $\widehat{\mathbb{N}^{\times}}$.
We can consider the logic $T_{\mathbb{Q}}$ which is the first order theory of all ordered subgroups of $\mathbb{Q}$ and $T_{\Delta(\mathbb{Q})}$ the first order theory of $Perf^{\mathbb{Q}}$. Then we have:
Theorem 6.3.
The first order theory $T_{\mathbb{Q}}$ is recursively axiomatizable.
Proof.
In [37] we find a complete characterization of a class of ordered abelian groups, called regularly ordered, which includes archimedean groups. We take from [37] the definitions of regularly ordered group and $Gp$.
If $G$ is a group, and $p$ is a prime number, $Gp$ is the largest finite number of elements of $G$ pairwise non-congruent modulo $p$, or $\infty$ if this number does not exist.
A group is discretely ordered if is has a smallest positive element.
A group is regularly discrete if $G$ is discretely ordered and such that $Gp=p$ for every prime $p$.
$G$ is regularly dense if, for any positive integer $p$ and any elements $a<b$, there is $x$ between $a$ and $b$ which is divisible by $p$.
$G$ is regularly ordered if it is regularly discrete or regularly dense.
The axioms of $T_{\mathbb{Q}}$ are: 1) $Gp\leq p$ for every prime $p$. 2) $G$ is regularly ordered.
Clearly 1) is expressible by (a recursive set of) first order axioms. For 2), note that $G$ is regularly ordered if it has an atom and for every pair of integers, $p,q$, either $G$ has an atom and $p$ is prime and is $p$-regularly discrete, or $G$ is $q$-regularly dense. So, also 2) is expressible by (a recursive set of) first order axioms.
In fact, by [37], every regularly ordered group with $Gp\leq p$ for every prime $p$ is elementarily equivalent to a subgroup of $\mathbb{Q}$.
Conversely, if $G$ verifies the theory of all the subgroups of $\mathbb{Q}$, then it is regularly ordered, and verifies $Gp\leq p$ for every $p$.
Theorem 6.4.
The class of ordered subgroups of $\mathbb{Q}$ is geometric in the language of ordered groups.
Proof.
$G\subseteq\mathbb{Q}$ if and only if $G$ is a group, $G$ is torsion free, and for every $x,y\not=0\in G$ there are $m,n\in\mathbb{N}$ such that $mx=ny$. Furthermore, denoting the order of $G$ by $<_{G}$, then the following conditions hold:
$\forall x$ $\forall y$ $\exists n$ $\exists m$ such that: $m,n\in Z$
  • $ny=mx$
  • (($na=mx$) and ($qb=px$)) implies (($a<_{G}b$ iff $mq<np$)), with $p,q\in\mathbb{Q}$.
All these properties are geometric. ∎
Theorem 6.5.
The class $Perf^{\mathbb{Q}}$ is geometric in the language of $MV$-algebras..
Proof.
Let $A\in Perf^{\mathbb{Q}}$. Then $(\Delta(A))^{+}\subseteq\mathbb{Q}$, that is equivalent to the axiom:
$\forall x$ $\forall y$ $\exists n$ $\exists m$ such that: $m,n\in Z^{+}$
  • $ny=mx$
  • (($na=mx$) and ($qb=px$)) implies (($a<_{G}b$ iff $mq<np$)), with $p,q\in\mathbb{Q}$.
All these properties are geometric. ∎
The previous results give information on the theory of certain subgroups of $\mathbb{Q}$. By using the functor $\Delta$ one can transfer this information on perfect MV-algebras.
Theorem 6.6.
The first order theory $T_{\mathbb{Q}}^{\Delta}$ of $Perf^{\mathbb{Q}}$ is mutually many one reducible with $T_{\mathbb{Q}}$. In particular it is recursively axiomatizable.
Proof.
The theorem follows from Theorem 6.2 and Theorem 5.2. ∎
The Theorem 6.2 has some corollaries.
Corollary 6.2.
Every two perfect MV-chains with divisible radical are elementarily equivalent.
Proof.
It is well known that every two divisible totally ordered abelian groups are elementarily equivalent. Moreover $G$ is divisible if and only if $\Delta(G)$ has divisible radical. Now the result follows from Theorem 6.2. ∎
Corollary 6.3.
The first order theory of perfect MV-chains is decidable.
Proof.
The theory of perfect MV-chains reduces to the theory of totally ordered abelian groups, which is decidable by [23]. ∎
The following result is a kind of converse of theorem 6.2, based on [10] prop 2.4.2.
We go from an abelian $\ell$-group $G$ to $G^{+}$.
Lemma 6.3.
Let $\phi$ be a formula in the language of lattice ordered groups. There is a formula $\Delta^{-1}+(\phi)$ in the language of lattice ordered monoids such that $G\models\phi(x_{1},\ldots,x_{n})$ if and only if $G^{+}\models\Delta^{-1}+(\phi)(x_{1}^{-},x_{1}^{+},\ldots,x_{n}^{-},x_{n}^{+}).$ Recall $x^{+}=x\vee 0$ and $x^{-}=x\wedge 0$.
Proof.
$x_{i}$ is replaced by $x_{i}^{-}-x_{i}^{+}$, and $\exists y.\psi(y)$ is replaced with $\exists z,w\in G^{+}.\psi(z-w)$. $\ell$-group Terms are inductively replaced by differences of lattice monoid terms as follows (see [10] prop 2.4.2 on how to construct the Grothendieck group of a lattice ordered monoid).
  • $0=0-$ 0 - 0;
  • $-(a-b)=b-a$;
  • $(a-b)+(c-d)=(a+c)-(b+d)$;
  • $(a-b)\wedge(c-d)=((a+d)\wedge(b+c))-(b+d)$;
  • $(a-b)\vee(c-d)=((a+d)\vee(b+c))-(b+d)$.
At the end every equality of $\ell$ group terms $t-u=v-w$ is replaced by $t+w=u+v$ which is a formula of lattice ordered monoids.
Now the lattice ordered monoids $(G^{+},+,\wedge,\vee)$ and $(Rad(\Delta(G)),\oplus,\wedge,\vee)$ are isomorphic. So we have:
Corollary 6.4.
Let $\phi$ be a formula in the language of lattice ordered groups. There is a formula $\Delta^{-1}(\phi)$ in the language of MV-algebras such that $G\models\phi(x_{1},\ldots,x_{n})$ if and only if $\Delta^{-1}(G)\models\Delta^{-1}(\phi)(x_{1}^{-},x_{1}^{+},\ldots,x_{n}^{-},x_{n}^{+}).$ Recall $x^{+}=x\vee 0$ and $x^{-}=x\wedge 0$.
We obtain the following model theoretic information:
Corollary 6.5.
The theory of perfect MV-algebras is undecidable.
Proof.
This follows by the previous lemma since the theory of all abelian $\ell$-groups is undecidable, see [5]. ∎
Finally we perform translations between first order formulas but we do not introduce any quantifier. So we can use these translations to prove model theoretic results like elimination of quantifiers or model completeness for theories of perfect MV-algebras.
We begin by observing:
Lemma 6.4.
The formula $\Delta(\phi)$ has the same quantifier depth as $\phi$ and , if $\phi$ is a positive boolean combination of atomic formulas, then so is $\Delta(\phi)$.
The same holds for the formula $\Delta^{-1}(\phi)$.
This gives, for instance, the following results: In terms of classical model theory we obtain:
Theorem 6.7.
The theory of perfect MV-chains with divisible radical is the model completion of the theory of perfect mv chains.
Proof.
This follows from the theorem of [28] that the theory of divisible MV chains is the model completion of the theory of mv chains. ∎
Theorem 6.8.
The theory of perfect MV-chains with divisible radical verifies the elimination of quantifiers. In particular, this theory is model complete, and every element of $Perf^{\mathbb{Q}}$ is an elementary substructure of $\Delta(\mathbb{Q})$.
Proof.
The theory T of divisible totally ordered abelian groups verifies the elimination of quantifiers. The perfect MV-algebra $\Delta(G)$ has a divisible radical if and only if $G$ is a divisible totally ordered abelian group. In $T$, every formula is equivalent to a quantifier free formula. By the previous subsection, the same holds for the theory of perfect MV-chains with divisible radical. ∎
Finally, using some of above results of this section, we obtain algebraic and model theoretic properties of the points of the topos $\widehat{\mathbb{N}^{\times}}$. We start with
Theorem 6.9.
([12], theorem 2.1) The category $\bf{Pt}$$(\widehat{\mathbb{N}^{\times}})$ of the points of the topos $\widehat{\mathbb{N}^{\times}}$ is equivalent to the category $Ab^{\mathbb{Q}}$ of Abelian $\ell$-groups included in $\mathbb{Q}$ where only increasing group homomorphisms are considered as morphisms.
By composing with the functor $\Delta$ we obtain:
Corollary 6.6.
The category $\bf{Pt}$$(\widehat{\mathbb{N}^{\times}})$ of the points of the topos $\widehat{\mathbb{N}^{\times}}$ is equivalent to the category $Perf^{\mathbb{Q}}$.
Let us introduce a further definition.
Definition 6.3.
If $\mathbf{p}\in Pt(\mathbb{N}^{\times}))$, let $Pt_{\mathbb{Q}}(\mathbf{p})$ be the corresponding subgroup of $\mathbb{Q}$. The map $pt_{\mathbb{Q}}$ is a functor. So for every $\mathbf{p}\in Pt(\mathbb{N}^{\times})$ we define
$\Theta_{pt}(\mathbf{p})=\Theta_{\mathbb{Q}}((\Delta^{\mathbb{Q}}(Pt_{\mathbb{Q}}(\mathbf{p})))$
and $\Theta_{pt}$ is a functor from $Pt(\widehat{\mathbb{N}^{\times}})$ to $C\ell BS$.
So we have
Corollary 6.7.
There is a natural functor from the category $Pt(\widehat{\mathbb{N}^{\times}})$ to the category Coupled $\ell$-Bisemirings.
Given the geometric characterization, via a categorical equivalence, of algebras from $Perf^{\mathbb{Q}}$ as points of the natural topos $\widehat{\mathbb{N}^{\times}}$, it would be interesting to state results about the first order theory and geometric theory of $Perf^{\mathbb{Q}}$ to get logical information about the points of $\widehat{\mathbb{N}^{\times}}$.
Indeed, we have a threefold interaction among the geometric side, points of the topos $\widehat{\mathbb{N}^{\times}}$, the algebraic side of ordered subgroups of $\mathbb{Q}$ and the logical side of $Perf^{\mathbb{Q}}$, models of an extension of $\L$ukasiewicz logic.
Geometric logic turns out to be the right logic for topos theory. Every geometric theory has a classifying topos, and when two theories have the same topos (that is they are Morita equivalent) they have many properties in common, despite they may look very different. For instance, the theories of MV-algebras and abelian $\ell$-groups with strong unit are Morita equivalent, see [7], and the theories of perfect MV-algebras and abelian $\ell$-groups are Morita equivalent, see [8]. So, it seems natural to look for a geometric logic of $\widehat{\mathbb{N}^{\times}}$.
Given the categorial equivalence between $Perf^{\mathbb{Q}}$ and $Pt(\widehat{\mathbb{N}^{\times}})$, we derive:
Corollary 6.8.
The first order theory of the points of the topos $\widehat{\mathbb{N}^{\times}}$ is recursively axiomatizable.
Proof.
A point $\mathbf{p}\in\widehat{\mathbb{N}^{\times}}$ is a set $X$ together with a $\mathbb{N}$-indexed family of functions $f_{n}$ from $X$ to $X$. Suppose $\sigma$ is a sentence of the language of $Pt(\widehat{\mathbb{N}^{\times}})$, and $\sigma$ is true in a point $\mathbf{p}\in Pt(\widehat{\mathbb{N}^{\times}})$ where $X$ is a subgroup $H$ of $\mathbb{Q}$ and $f_{n}(x)=nx$. Then $\sigma$ is true in some subgroup of $\mathbb{Q}$. Likewise, if $\sigma$ is true in a subgroup of $\mathbb{Q}$, then $\sigma$ is true in point of $\widehat{\mathbb{N}^{\times}}$.
So, the first order theory of $Pt(\widehat{\mathbb{N}^{\times}})$ is the same of that of the subgroups of $\mathbb{Q}$ and is recursively axiomatizable.
6.1 Flatness and $Pt(\widehat{\mathbb{N}^{\times}})$
In [12] theorem 2.1 holds because the points of $\widehat{\mathbb{N}^{\times}}$ correspond to flat functors from $\mathbb{N}^{\times}$ to Set, we can call them flat actions of $\mathbb{N}^{\times}$ and we denote by $Flact(\mathbb{N}^{\times})$ their category.
Definition 6.4.
Recall that a functor $F:\mathbb{N}^{\times}\to Set$ is flat if
  1. 1.
    $F(*)=X\not=\emptyset$, where * is the unique object of $\mathbb{N}^{\times}$
  2. 2.
    given $y,z\in X$ there are $m,n\in\mathbb{N}$ and $w\in X$ such that $mw=y$ and $nw=z$
  3. 3.
    if $m,n\in\mathbb{N}$ and $y\in X$ such that $my=ny=w$, then there is $p\in\mathbb{N}$ and $z\in X$ such that $mp=np$ and $pz=y$.
Note that all flatness axioms above are expressed in the geometric logic in the sense of topos theory. In particular, flat functors are closed under colimits.
In lemma 2.2 of [12] we find:
Lemma 6.5.
There is a categorical equivalence between the category $Flact(X)$ and the category of ordered subgroups of the rationals.
Proof.
Given an action on $X$, the idea is to define a sum on $X$ by $kz+k^{\prime}z=(k+k^{\prime})z$. This is well defined since the action is flat, and gives a semigroup which is the positive part of a subgroup of the rationals. Conversely, the Frobenius action on the positive part of a subgroup of the rationals is flat. ∎
We remark that the point $\mathbf{p_{\mathbb{Q}}}\in\widehat{\mathbb{N}^{\times}}$ corresponding to $\mathbb{Q}$ is the exponentiation functor $F_{\mathbb{Q}}:\mathbb{N}\times X\to X$, with $X=\mathbb{Q}^{+}$ and $q\in\mathbb{Q}^{+}$, $F_{\mathbb{Q}}(n,q)\rightarrow q^{n}$.
Let $L_{Pt}$ be the language with a sequence of functional symbols $\{F_{n}:\mathbb{N}\times X\rightarrow X\}$. Then we have:
Theorem 6.10.
The class $Pt(\widehat{\mathbb{N}^{\times}})$ is geometric over the language $L_{Pt}$.
Corollary 6.9.
Every point $\mathbf{p}\in Pt(\widehat{\mathbb{N}^{\times}})$ is an elementary substructure of $\mathbf{p_{\mathbb{Q}}}$.
7 Conclusions
As stated in the Introduction, the purpose of this article is to illustrate the interactions between algebraic structures such as Semirings, Tropical structures and $\L$ukasiewicz logic.
This perspective has led to a generalisation of the concept of Tropicalisation of algebraic structures in a form that opened up an unexpected interpretation of an extension of $\L$ukasiewicz logic. This logic exhibits models that are functorially connected to a category of points of a non-commutative geometry.
Thus, assuming the functorial version of the tropicalization of an algebraic structure, the above interation, among others, takes form as follows:
  • from the algebraic side arise the new algebraic categories of Lattice Bisemirings and of Coupled Lattice Bisemirings,
  • from the geometric side, the category of points of the Topos $\widehat{\mathbb{N}^{\times}}$ results equivalent to several algebraic categories,
  • from the logical side we have that the theory of points of $\widehat{\mathbb{N}^{\times}}$ is recursively axiomatizable, and the theory of flat actions is geometric.
As a final comment we can say that the Tropicalization of Universal Geometry of $\L$ukasiewicz logics ([35], [3]) seems to be a rich topic which deserves to be addressed. This means extending the perspective of the interactions exhibited in the present article.